OpenStudy (anonymous):

Arithmetic means are inserted between 1 and 50 so that the sum of the resulting arithmetic progression is 561. (a) How many arithmetic means were inserted? (b) What is the common di fference of the resulting arithmetic progression?

OpenStudy (anonymous):

here clearly the first term is 1 and the last term is 50 thes sum of AP whose first and last term is given= (n/2)*(first term +last term)

OpenStudy (anonymous):

hence the number of term=2*561/(1+50) =2*561/51=2*11=22 50=1+(11-1)*d d is the common difference

OpenStudy (anonymous):

Isn't it 50=1+(22-1)d?

OpenStudy (anonymous):

THANKYOOOU. :)

OpenStudy (jack1):

wait... was the answer as @puuhreey said: 50=1+(22-1)d so d = 49/21 = 2.333 or was it as @matricked said and : 50=1+(11-1)*d so d = 49/10 = 4.9 ...?